Redshift measurements show that a galaxy is moving away from the Earth at the speed of 60000 km/s. Taking the Hubble constant H, as 2.2 x 10-18 s-1 and 1 light-year as 1013 km, calculate and show your working for: a how far the galaxy is from the Earth


1
Expert's answer
2022-05-30T15:17:46-0400

"r=\frac vH=27.3\cdot 10^{21}~km=27.3\cdot 10^8~l.y."